An investment project provides cash inflows of $680 per year for 10 years.
a. What is the project payback period if the initial cost is $2,040?

  

b. What is the project payback period if the initial cost is $3,808?

  

c. What is the project payback period if the initial cost is $7,480?
Answer #1

payback period=initial cost/annual cash flows

1.payback period=2040/680=3 years.

2.payback period=3808/680=5.6 years

3.payback period=7480/680=Never